(-3ab)2 (-2ab)3
=(-3 to 2nd a2 b2)(-2 to 3rd a3 b3)
=9*8 a2+3 b2+3
=-72 a5 b5

wasn't quite sure how to type out line 2
but I'm tyring to figure out how to determine whether or not an answer is + or -

Answer by tutorcecilia(2144)

(-3ab)^2 (-2ab)^3 [Use CAP 6 key for the exponent sign]
((-3)^2a^2b^2) ((-2)^3a^3b^3) [I added parenthesis for emphasis]
(9a^2b^2)(-8a^3b^3)
(9)(-8)a^5b^5
-72a^5b^5
Otherwise, I agree with your answer. You did a good job!!!
Here are some rules to remember:
-3^2 = (-1)3^2 = (-1)(9) = -9 [The negative sign does not belong to the 3, it is really multiplying by (-1)]
.
(-3)^3 = (-3)(-3) = 9 [The negative sign belongs to the 3 because it is within the parenthesis]
.
-3^3 = (-1)(3)^3 = (-1)(27) = -27
.
(-3)^3 = (-3)(-3)(-3) = (9)(-3)= -27